# Retrieve to Explain: Evidence-driven Predictions with Language Models

### v2

@@ Abstract
Machine learning models, particularly language models, are notoriously difficult to introspect. Black-box models can mask both issues in model training and harmful biases. For human-in-the-loop processes, opaque predictions can drive lack of trust, limiting a model’s impact even when it performs effectively. To address these issues, we introduce Retrieve to Explain (R2E). R2E is a retrieval-based language model that prioritizes amongst a pre-defined set of possible answers to a research question based on the evidence in a document corpus, using Shapley values to identify the relative importance of pieces of evidence to the final prediction. R2E can adapt to new evidence without retraining, and incorporate structured data through templating into natural language. We assess on the use case of drug target identification from published scientific literature, where we show that the model outperforms an industry-standard genetics-based approach on predicting clinical trial outcomes.
Retrieve to Explain, R2E, Machine learning, Explainability, Retrieval, Language model, Shapley values, Human-in-the-loop, Multimodal, Drug discovery, Target identification, Genetics, Biomedical
@@ 1 Introduction
Recent research has shown that language models can act as knowledge bases, supplying answers to factual user queries on-demand using only the learned parameters (Petroni et al., 2019; Brayne et al., 2022). Language models can also be supplied with searchable knowledge bases for a retrieval-augmented approach to question answering (Chen et al., 2017; Lewis et al., 2020; Izacard & Grave, 2021).
@@
Figure 1: R2E drug target identification example. R2E makes predictions based on retrieved evidence and provides explanations in terms of the evidence. Query: User queries are phrased in cloze-style, where [MASK] can be filled from a set of answers (named entities). Retrieval: R2E retrieves the evidence most relevant to the query for each possible answer. Prediction: The model generates a score for each answer based on the evidence. Explanation: Predictions are explained using Shapley values of the evidence. Here, the best evidence is indirect, based on the role of CD6 in mechanisms central to rheumatoid arthritis pathology.
@@
Beyond answering factual queries, a searchable knowledge base could provide evidence for queries without known answers, including scientific research questions (e.g. What are some promising drug targets to treat rheumatoid arthritis?). By proposing new hypotheses supported by both direct and indirect scientific evidence, AI models could support and accelerate the scientific discovery process (Paliwal et al., 2020; Aliper et al., 2023; Sourati & Evans, 2023).
@@
When acting on model hypotheses is costly or risky, an explainable model can mitigate risk by allowing a human expert to inspect the evidence and reasoning behind each prediction before acting on it (i.e. a human-in-the-loop setup). Explainability can also help to identify model flaws or systemic biases, leading to improved performance and task alignment (Kulesza et al., 2015).
@@
Here, we introduce Retrieve to Explain (R2E), an approach for explainable language model predictions (Figure 1). Given a cloze-style user query, R2E first retrieves the most relevant evidence from an evidence corpus, partitioned according to each possible answer; we consider a set of answers comprised of named entities. The model then scores each answer based on its supporting evidence to generate a fully ranked list. The R2E architecture is designed to ensure that potential answers are represented explicitly in terms of their supporting evidence. In particular, the feature space is comprised of the evidence itself, allowing the use of feature attribution methods to estimate the contribution of each piece of evidence to the final prediction; here, we use Shapley values (Shapley et al., 1953; Lundberg & Lee, 2017). In addition to clarifying the prediction to the user, we show that this evidence-oriented approach allows model predictions to be updated without retraining by modifying the corpus, such as updating it with new evidence. Since R2E can generate a score for every answer in the answer set, it is particularly applicable in human-in-the-loop scenarios where many potential hypotheses should be prioritized for user review. In addition, this approach allows users to investigate negative, low scoring predictions.
@@
With half of drugs failing to show efficacy when tested in human populations (Wong et al., 2019), we focus here on drug target identification, a critical early stage in drug discovery where specific genes or proteins (targets) are selected as the focus for developing treatments. For this task, we train R2E to score protein-coding genes by relevance to a user query based on a scientific literature corpus. We then augment the corpus with genetic associations by templating them into natural language, allowing the model to use both evidence sources. We show how Shapley values on the evidence correlate with Large Language Model (LLM) relevance assessments, which similarly correlate with human experts. Notably, when used to predict clinical trial outcomes, R2E significantly outperforms genetics evidence, a widely recognised predictor in the pharmaceutical industry (Nelson et al., 2015; Trajanoska et al., 2023). R2E outperforms the genetics baseline even when supplied only with genetics evidence, suggesting that representing gene-disease associations in natural language improves generalization. Finally, we show that explainability enables the use of LLMs to audit prediction reasoning, further improving performance.
@@ 2.1 Language Models with Retrieval
Many language models have been proposed to leverage retrieved training examples at inference time (Khandelwal et al., 2019; Karpukhin et al., 2020; Guu et al., 2020; Lewis et al., 2020; Yogatama et al., 2021; Izacard & Grave, 2021; Borgeaud et al., 2022; Izacard et al., 2022). R2E differs from these existing approaches by (1) scoring all possible answers in a given answer set and (2) attributing each answer’s score to evidence passages using Shapley values. This approach follows from the application: R2E is designed for answering research questions that merit deep user engagement (e.g. identifying drug targets for a disease) as opposed to typical factual recall tasks (e.g. identifying a country’s capital city). In the former case, there is a premium on generating multiple possible answers and explanations, allowing a human user to investigate inference results in depth.
Of the existing architectures, most are sequence-to-sequence models; this approach makes it intractable to score (or even enumerate) the entire answer set. R2E perhaps bears the most resemblance to kNN-LM (Khandelwal et al., 2019) which uses retrieval to improve next-token prediction. However, kNN-LM uses retrieval to augment a standard masked language model, while R2E is fully retrieval-based to enable evidence-driven explanations. The Fusion-in-Decoder (FiD) approach (Izacard & Grave, 2021) also bears a notable resemblance to R2E; both merge each piece of evidence with the query independently before jointly processing evidence. While FiD is motivated by efficiency and performance, we are additionally motivated here by explainability.
2.2 Explainability
R2E is inspired by SHAP (SHapley Additive exPlanations) (Lundberg & Lee, 2017), which explains model predictions by approximating Shapley values (Shapley et al., 1953). SHAP evaluates how input features impact model predictions; in part, by hiding subsets of features. Here, we are instead interested in how evidence drives the model prediction. While Data Shapley and related methods (Ghorbani & Zou, 2019; Pandl et al., 2021) have examined the Shapley values of training data, they focus on the impact of training examples on overall model performance rather than inference-time predictions. In fact, evaluating data Shapley values for predictions is typically impossible with fully parametric approaches; after training data has been compressed into the model parameters, there is no straightforward way to hide it for an arbitrary prediction. R2E instead uses a semi-parametric approach where evidence contributes to the prediction at inference time; by turning evidence into features, we can hide evidence and re-evaluate the prediction.
Alternatively, SimplEx (Crabbé et al., 2021) explains predictions by approximating a classified input in terms of a corpus of classified exemplars. SimplEx is general-purpose but indirect: the corpus illuminates black-box model predictions, but does not impact them. In contrast, the corpus directly drives model predictions in R2E, and changing the corpus will change predictions.
@@ 2.3 Models for Hypothesis Generation
The use of models in generating or evaluating scientific hypotheses is an emerging area of research. Knowledge graphs are a popular approach for novel hypothesis generation, because their structure enables multi-hop inference between unconnected nodes. Novel hypotheses have been generated by subject-area experts directly querying and inspecting a knowledge graph (Smith et al., 2021).
Sourati & Evans (2023) use knowledge graph patterns for material property prediction and drug re-purposing, additionally leveraging nodes for specific researchers to infer which discoveries are more or less likely to be discovered based on social dynamics. Paliwal et al. (2020) used tensor factorization on a biomedical knowledge graph to predict future research findings and clinical trial outcomes for therapeutic drug targets. Aliper et al. (2023) similarly employed a biomedical knowledge graph to predict clinical trial outcomes; they used a graph transformer network, and ensembled it with a tabular model leveraging clinical trial design features. R2E differs from these approaches by enabling explainability of predictions from the evidence and operating directly on published research without the need to construct a knowledge graph.
In this vein, Tshitoyan et al. (2019) work with a materials science research corpus to identify new material properties. They use cosine similarity on unsupervised word embeddings, specifically word2vec (Mikolov et al., 2013). This approach resembles our parametric masked language model baseline, except that in our case embeddings are derived using a transformer. Tshitoyan et al. suggest that word2vec enables indirect inference similar to that in a knowledge graph; for instance, a material that has never been defined as thermoelectric may nonetheless be mentioned alongside properties associated with thermoelectricity. We observe a similar phenomenon in R2E: for instance, a target never identified directly with a disease may still have been shown to regulate disease-relevant mechanisms (Figure 1) or to be genetically associated with relevant traits (Appendix T). R2E can use these indirect findings as support.
@@ 3 Methods
We consider the problem of scoring $N$ potential answers $\mathcal{A}=\{a_{i}\}_{1}^{N}$ to a user query $q$ such that they can be ranked from most to least relevant. We restrict $q$ to be a cloze-style statement as opposed to a natural question (e.g. [MASK] is a promising drug target for the treatment of osteoporosis.), where each answer $a_{i}$ represents a potential named entity for the mask location. Lewis et al. (2019) provides an approach to translate between cloze- and natural-style questions.
Figure 2: R2E architecture schematic. Illustration of R2E inference and explanation. Input: A user-defined cloze-style query, an entity to evaluate, and a corpus of evidence passages corresponding to that entity with entity mentions replaced with [MASK]. Retriever: The query text is encoded with a transformer. All of the entity’s evidence passages are encoded prior to inference, using the same encoder, and stored in a FAISS search index. The $k$ evidence passages with highest cosine similarity to the query are retrieved. Reasoner: The query/evidence pair embeddings are layer-normalised. Each pair is combined at corresponding dimensions into a single embedding using convolutional layers. All combined pair embeddings are passed to a set transformer, followed by a linear layer and sigmoid to obtain the binary probability. Shapley values for each pair (corresponding to each piece of evidence) are computed. Output: To rank over a set of entities, binary probabilities are obtained independently for each entity as described above. Shapley values attribute model predictions back to the evidence passages providing an explanation of the model’s prediction.
@@ 3.1 Masked Entity-Linked Corpus
Our approach uses a training corpus of textual passages, $\mathcal{D}$, each containing at least one named entity from the answer set $\mathcal{A}$. An entity linking methodology identifies named entities in $\mathcal{A}$ in the corpus.
For each passage, the span of every occurrence of the named entity is replaced by a [MASK] token. When multiple unique answers occur, we generate a new passage with each masked in turn while the others appear as plain text. Each example is therefore a tuple $(a,d)$ consisting of an entity identifier and a masked text passage in which that entity occurs.
In our specific use case of drug target identification, $\mathcal{A}$ consisted of 19,176 protein-coding gene entities, hereafter referred to collectively as Genes. Surface forms of the gene and corresponding protein (encoded by the gene), were grounded to the same entity under a 1:1 assumption. $\mathcal{D}$ was a large-scale corpus of 160 million sentences from scientific literature. For more details on the corpus and splits used, including temporal splits, see Appendix B; for details on the entity linking methodology, see Appendix C. Note that $|\mathcal{D}|$ could support other use cases (e.g. biomarker identification, drug repurposing, biological mechanism selection) by adjusting the named entity answer set.
@@ 3.2 Masked Language Model (MLM)
We first consider a parametric approach to the problem based on the pre-training method in Brayne et al. (2022). We use an encoder-only transformer (Vaswani et al., 2017), specifically a scaled-down version of PubMedBERT (Gu et al., 2021). For query passages $d^{q}\in\mathcal{D}$ containing a masked answer $a_{i}\in\mathcal{A}$, we train the model to predict $p(a_{i}|d^{q})$.
The query embedding is the mean over output embeddings corresponding to [MASK] tokens when the entity is mentioned multiple times. We take the dot product with a learned embedding for each possible answer in the answer set, then apply a bias and softmax to predict $p(a_{i}|d^{q})\,\forall\,a_{i}\in\mathcal{A}$. We train with a cross-entropy loss. Pre-trained weights for domain-specific models are available (e.g. PubMedBERT), but we train from scratch to avoid leakage from pre-training data in our literature evaluations.
This model is both a baseline (MLM) and the basis for the Retriever component of R2E (Section 3.3).

### v4

# Retrieve to Explain: Evidence-driven Predictions for Explainable Drug Target Identification

@@ Abstract
Language models hold incredible promise for enabling scientific discovery by synthesizing massive research corpora. Many complex scientific research questions have multiple plausible answers, each supported by evidence of varying strength. However, existing language models lack the capability to quantitatively and faithfully compare answer plausibility in terms of supporting evidence. To address this, we introduce Retrieve to Explain (R2E), a retrieval-based model that scores and ranks all possible answers to a research question based on evidence retrieved from a document corpus. The architecture represents each answer only in terms of its supporting evidence, with the answer itself masked. This allows us to extend feature attribution methods such as Shapley values, to transparently attribute answer scores to supporting evidence at inference time. The architecture also allows incorporation of new evidence without retraining, including non-textual data modalities templated into natural language. We developed R2E for the challenging scientific discovery task of drug target identification, a human-in-the-loop process where failures are extremely costly and explainability paramount. When predicting whether drug targets will subsequently be confirmed as efficacious in clinical trials, R2E not only matches non-explainable literature-based models but also surpasses a genetics-based target identification approach used throughout the pharmaceutical industry.
Retrieve to Explain: Evidence-driven Predictions for Explainable Drug Target Identification
Ravi Patel   Angus Brayne   Rogier Hintzen Daniel Jaroslawicz   Georgiana Neculae   Dane Corneil BenevolentAI, London, United Kindgom {ravi.patel, dane.corneil}@benevolent.ai
@@ 1 Introduction
Language models can act as knowledge bases, supplying answers to factual user queries using only the learned parameters Petroni et al. (2019); Brayne et al. (2022). They can also be provided with access to searchable knowledge bases for retrieval-augmented question answering Chen et al. (2017); Lewis et al. (2020); Izacard and Grave (2021).
@@
Figure 1: R2E drug target identification example. R2E makes predictions based on retrieved evidence and provides explanations in terms of the evidence. Query: User queries are phrased in cloze-style, where [MASK] can be filled from a set of potential answers (named entities). For target identification, answers are the set of protein-coding genes (potential drug targets), and the query specifies a disease. Retrieval: R2E retrieves the evidence most relevant to the query for each potential answer, where evidence here is taken from across the biomedical literature that mentions the specific answer. Prediction: The model scores each answer based on the supporting evidence. Explanation: Each answer score is directly and quantitatively attributed to its retrieved evidence using Shapley values. Here, the best evidence is indirect, based on the role of CD6 in mechanisms central to rheumatoid arthritis pathology.
@@
Beyond answering factual queries, a searchable knowledge base could provide evidence for queries without known answers, including scientific research questions (e.g. What are some promising drug targets to treat rheumatoid arthritis?). By proposing new hypotheses supported by both direct and indirect scientific evidence, AI models could facilitate scientific discovery Paliwal et al. (2020); Aliper et al. (2023); Sourati and Evans (2023).
@@
For high-stakes settings where acting on model hypotheses is costly or risky, an explainable model can mitigate risk by allowing a human expert to inspect the evidence and reasoning behind each prediction before acting on it (human-in-the-loop). Explainability can also help to identify model flaws or systemic biases, leading to improved performance and task alignment Kulesza et al. (2015).
@@
Here, we introduce Retrieve to Explain (R2E), an approach for language model prediction with faithful and quantitative explanations (Figure 1). Given a cloze-style user query, R2E first retrieves the most relevant evidence from an evidence corpus, partitioned according to each possible answer. We consider a set of answers comprised of named entities. The model then scores each answer based on its supporting evidence to generate a ranked list. The R2E architecture represents potential answers explicitly in terms of their supporting evidence. In particular, the feature space is the evidence itself, enabling explainability with feature attribution methods to infer the contribution of each piece of evidence to the prediction. Here, we use Shapley values Shapley et al. (1953); Lundberg and Lee (2017). In addition to explainability, we show that this evidence-oriented approach allows model predictions to be updated without retraining by modifying the corpus, such as introducing new evidence. Since R2E can generate a score for every answer in the answer set, it is particularly applicable in human-in-the-loop scenarios where many potential hypotheses are prioritized for user review.
@@
With half of drugs failing to show efficacy when tested in human populations Wong et al. (2019), often due to an ineffective choice of drug target, we developed R2E for drug target identification. Target identification is an especially challenging scientific discovery problem where specific genes or proteins (targets) are selected as the focus for developing new treatments, and where failures are extremely costly Olivier J. Wouters (2020). We train R2E to score protein-coding genes by relevance to a user query based on a scientific literature corpus. We then augment the corpus with genetic associations by templating them into natural language, allowing the model to use both evidence sources. We show that Shapley values on individual pieces of evidence correlate with large language model (LLM) relevance assessments, which similarly correlate with human experts. Notably, when used to predict clinical trial outcomes, R2E significantly outperforms both genetics evidence, a widely recognised predictor in the pharmaceutical industry Nelson et al. (2015); Trajanoska et al. (2023), and a few-shot, chain-of-thought, retrieval-augmentation GPT-4 baseline, a setup that in practice would also be prohibitively costly and sacrifices faithful explainability. R2E outperforms the genetics baseline even when supplied only with genetics evidence, suggesting that representing gene-trait associations in natural language improves generalization over a structured ontology. Finally, we show that R2E’s explainability enables the use of LLMs to audit prediction reasoning, further improving performance.
@@
Our core contributions are as follows:
  • We introduce R2E, a novel architecture for retrieval-based high-stakes question answering, which scores the plausibility of each answer directly in terms of its supporting evidence, and thereby enables faithful, quantitative explainability using evidence-level Shapley values.
  • We develop R2E for the challenging scientific discovery problem of drug target identification; it is not only as predictive of clinical trial outcomes as non-explainable literature-based baselines, but also surpasses a genetics approach used throughout the pharmaceutical industry.
  • We release three new benchmarks to address the lack of publicly-available datasets for drug target identification and drive progress on this important scientific discovery problem.
@@ 2.1 Language Models with Retrieval
Many language models leverage retrieved text at inference time for question answering Khandelwal et al. (2019); Karpukhin et al. (2020); Guu et al. (2020); Lewis et al. (2020); Lee et al. (2020); Izacard and Grave (2021); Borgeaud et al. (2022); Izacard et al. (2022). R2E differs from these existing approaches by (1) scoring all possible answers in an answer set and (2) faithfully and quantitatively attributing each answer’s score to evidence passages using Shapley values. This approach follows from the application: R2E is designed for answering research questions that merit deep user engagement (e.g. identifying potential drug targets for a disease) as opposed to typical factual recall tasks (e.g. identifying a country’s capital city). Scoring many possible answers with faithful explanations allows a human to investigate them.
R2E perhaps bears the most resemblance to kNN-LM Khandelwal et al. (2019) which uses retrieval to improve next-token prediction. However, kNN-LM uses retrieval to augment a standard masked language model, while R2E is fully retrieval-based to enable evidence-driven explanations. The Fusion-in-Decoder (FiD) approach Izacard and Grave (2021) also bears a resemblance to R2E; both merge each piece of evidence with the query independently before jointly processing. FiD is motivated by efficiency and performance. We are additionally motivated by explainability. As discussed in depth in Appendix V, faithfully explainable multi-label prediction with existing generative LLM architectures is largely infeasible.
2.2 Explainability & Data Attribution
R2E is inspired by SHAP (SHapley Additive exPlanations) Lundberg and Lee (2017), which explains model predictions by approximating feature-level Shapley values Shapley et al. (1953). R2E extends feature attribution methods like SHAP to data, by using a retrieval-based architecture in which the feature space is comprised of evidence. R2E therefore also contrasts with explainability-focused training data attribution (TDA) methods Hammoudeh and Lowd (2024), such as representer point selection Sui et al. (2021), which evaluates the impact of training examples on predictions. Instead, R2E uses the evidence in the corpus at inference time for both prediction and explanation. Among TDA methods, Data Shapley Ghorbani and Zou (2019) also assigns Shapley values to data. Data Shapley focuses on explaining model performance rather than inference-time predictions.
SimplEx Crabbé et al. (2021) explains predictions by approximating an input in terms of a corpus of classified exemplars. SimplEx is general-purpose but indirect: the corpus illuminates black-box predictions, but does not impact them. In contrast, the corpus drives model predictions in R2E.
@@ 2.3 Models for Hypothesis Generation
The use of models in generating or evaluating scientific hypotheses is an emerging area of research. Knowledge graphs (KGs) are a popular approach for novel hypothesis generation, because their structure enables multi-hop inference between unconnected nodes. Novel hypotheses have been generated by subject-area experts directly querying and inspecting a KG Smith et al. (2021).
Sourati and Evans (2023) use KG patterns for material property prediction and drug re-purposing, additionally leveraging nodes for specific researchers to infer which discoveries are more or less likely to be discovered based on social dynamics. Paliwal et al. (2020) used tensor factorization on a biomedical KG to predict future research findings and clinical trial outcomes for therapeutic drug targets. Aliper et al. (2023) similarly employed a biomedical KG to predict clinical trial outcomes; they used a graph transformer network ensembled with a tabular model leveraging clinical trial design features. R2E differs from these approaches by enabling explainability in terms of the evidence and operating directly on published research without needing to construct a KG.
In this vein, Tshitoyan et al. (2019) work with a materials science research corpus to identify new material properties. They use cosine similarity on unsupervised word embeddings, specifically word2vec Mikolov et al. (2013). This resembles our parametric masked language model baseline, except that in our case embeddings are derived using a transformer. Tshitoyan et al. suggest that word2vec enables indirect inference similar to that in a KG; for instance, a material never defined as thermoelectric may be mentioned alongside properties associated with thermoelectricity. We observe a similar phenomenon in R2E: for instance, a target never identified directly with a disease may still have been shown to regulate disease-relevant mechanisms (Figure 1) or to be genetically associated with relevant traits (Appendix U.7). R2E can use these indirect findings as support.
@@ 3 Methods
We consider the problem of scoring $N$ potential answers $\mathcal{A}=\{a_{i}\}_{i=1}^{N}$ to a user query $q$, to rank them from most to least relevant. To align with the training corpus (Section 3.1), we let $q$ be cloze-style (e.g. [MASK] is a promising drug target for the treatment of osteoporosis.), where each answer $a_{i}$ represents a potential named entity at [MASK]. Lewis et al. (2019) provides an approach to translate between cloze- and natural-style questions.
Figure 2: R2E architecture schematic. Illustration of R2E inference and explanation. Input: A user-defined cloze-style query, a possible answer (named entity) to evaluate, and a corpus of evidence passages corresponding to that answer entity with entity mentions replaced with [MASK]. Retriever: The query text is encoded with a transformer. All of the entity’s evidence passages are encoded prior to inference, using the same encoder, and stored in a FAISS search index. The $k$ evidence passages with highest cosine similarity to the query are retrieved. Reasoner: Each evidence embedding is stacked with the query embedding. The resulting query-evidence pairs are layer-normalised before each pair is combined at corresponding dimensions into a single embedding using convolutional layers. All combined pair embeddings are passed to a set transformer, followed by a linear layer and sigmoid to obtain the binary probability. Shapley values for each pair (corresponding to each piece of evidence) can be computed to quantitatively explain the prediction. Output: To rank a set of answer entities $a_{1...N}$, binary probabilities are obtained independently for each. Shapley values attribute model predictions back to the evidence passages providing an explanation of the model’s prediction.
@@ 3.1 Masked Entity-Linked Corpus
Our approach uses a training corpus of textual passages, $\mathcal{D}$, each containing at least one named entity from the set of answer entities $\mathcal{A}$. Entity linking identifies and grounds entities in $\mathcal{A}$ in the corpus. For each passage, the span of every occurrence of a single entity is replaced by a [MASK] token. When the passage contains multiple unique entities in $\mathcal{A}$, we duplicate the passage with each masked in turn while the others appear as plain text. Each example is therefore a tuple $(a,d)$ consisting of an answer entity identifier $a\in\mathcal{A}$ and a masked text passage $d\in\mathcal{D}$ in which that entity occurs.
In application to drug target identification, $\mathcal{A}$ consisted of 19,176 protein-coding gene entities, hereafter referred to collectively as Genes. $\mathcal{D}$ was an entity-linked corpus of 160 million sentences from scientific literature. For more details on the corpus and splits, including temporal splits to avoid leakage, see Appendix B; for entity linking see Appendix C. $\mathcal{D}$ could in theory support other tasks (e.g. biomarker identification, drug repurposing, biological mechanism selection) by adjusting $\mathcal{A}$.
@@ 3.2 Masked Language Model (MLM)
We first consider a parametric approach based on the pre-training method in Brayne et al. (2022). We use an encoder-only transformer Vaswani et al. (2017), specifically a scaled-down version of PubMedBERT Gu et al. (2021). For query passages $d^{q}\in\mathcal{D}^{\mathrm{MLM}}\subset\mathcal{D}$ containing a masked answer $a_{i}\in\mathcal{A}$, we train to predict $p(a_{i}|d^{q})$.
The query embedding is the mean over output embeddings at [MASK] token positions. We take the dot product with a learned embedding for each possible answer $a_{i}\in\mathcal{A}$, then apply a bias and softmax to predict $p(a_{i}|d^{q})\,\forall\,a_{i}\in\mathcal{A}$. We train with cross-entropy loss. Pre-trained domain-specific model weights are available (e.g. PubMedBERT), but we train from scratch to avoid leakage from pre-training data in our temporally-split evaluations.
This model is both a baseline (MLM) and the basis for the Retriever of R2E (Section 3.3).